OpenStudy (anonymous):

A swimming pool is being filled at the rate of 12 cubic yards per minute. If the pool is 18 yards long, 10 yards wide, and 3 yards deep, how many minutes will it take to fill the pool? A 45 minutes B 1,233 minutes C 101 minutes D 540 minutes

OpenStudy (kira_yamato):

\[\frac{dV}{dt} = 12\] V = 18(10)(3) = 540 t = 540/12 = 45

OpenStudy (anonymous):

Thanks.